# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems: A Computational Examination of the Korteweg de-Vries and Lorenz equations

### v1

@@ 6. Conclusion
Based on the results of the simulation presented in Section 4, we conclude that the AOT data assimilation algorithm is unsuitable for non-dissipative or partially dissipative dynamical systems, such as the KdV equation, Equation 2.6, and the partially dissipative Lorenz system, Equation 5.1. In fact, we conclude that the pathologic counterexamples used in Section 4.1 present an insurmountable challenge for any method of data assimilation to recover. In particular, we note that counterexample we utilize in Figure 1 will fail for any method of data assimilation, as the true solution has no discernible impact on the observed modes, and so it is impossible to reliably tell the true solution apart from the zero solution (or one of an infinite number of solutions one can readily construct) based solely on the observations. Moreover, the counter examples considered in Section 4.1 are not specific to the KdV equation. The particular solutions considered in 2.0.5 were constructed due to the periodic invariance induced by the convolutional Fourier mode structure of the nonlinear term (see 2.0.4), and thus one can construct similar counter examples for other non-dissipative systems with this same nonlinearity, such as the inviscid Burgers equations.
In addition to showing that data assimilation simply doesn’t work for certain pathological counterexamples, we note that the extent to which data assimilation does work for the classical KdV equation may be highly dependent on the choice of initial data for the nudged system, $v^{in}$. Indeed, while it was proved in [4] that for 2D NSE the choice of initial data $v^{in}$ was arbitrary, we have here that data assimilation will simply fail if not initialized properly, as it does in Figure 1. The problem of initialization is one of the core strengths of the nudging algorithm, and so it is worth noting that this property fails to hold for a system without finitely many determining modes.

### v2

@@ 5. Conclusion
Based on the results of the simulation presented in Section˜3.2 and Section˜4.2, we conclude that the AOT data assimilation algorithm is unsuitable for non-dissipative or partially dissipative dynamical systems, such as the incompressible Euler and KdV equations, Sections˜4.1 and 19, respectively, and the partially dissipative Lorenz system, Equation˜1. In fact, we conclude that the pathologic counterexamples used in Sections˜3.2.1 and 4 present an insurmountable challenge for any method of data assimilation to recover. In particular, we note that counterexample we utilize in Figure˜4 will fail for any method of data assimilation, as the true solution has no discernible impact on the observed modes, and so it is impossible to reliably tell the true solution apart from the zero solution (or one of an infinite number of solutions one can readily construct) based solely on the observations. Moreover, the counter examples we considered in this work are not specific to the KdV or Euler equations. The particular solutions considered in Proposition˜3.5 were constructed due to the periodic invariance induced by the convolutional Fourier mode structure of the nonlinear term (see Corollary˜3.4), and thus one can construct similar counter examples for other non-dissipative systems with a similar nonlinearity.
In addition to showing that data assimilation simply doesn’t work for certain pathological counterexamples, we note that the extent to which data assimilation does work for the classical KdV equation may be highly dependent on the choice of initial data for the nudged system, $v^{in}$. Indeed, while it was proved in [4] that for 2D NSE the choice of initial data $v^{in}$ was arbitrary, we have seen here that data assimilation will simply fail if not initialized properly, as it does in Figure˜4. The problem of initialization is one of the core strengths of the nudging algorithm, and so it is worth noting that this property fails to hold for a system without finitely many determining modes.